OpenStudy (waheguru):

How Do I solve for x 3 = -2x^2+x+3

OpenStudy (abb0t):

What two numbers when multiplied give you -6, but when added give you 1?

OpenStudy (waheguru):

-3 and 2

OpenStudy (abb0t):

Oh wait, i see there's a three. Subtract 3 from both sides and set it equal to zero. Then factor out an "x".

OpenStudy (waheguru):

yea i have 0 = 2x^2+x

OpenStudy (abb0t):

Factor out an "x".

OpenStudy (waheguru):

0= -x(2x-x)

OpenStudy (abb0t):

Not quite. When you factor out an "x" from the second term you shouldn't have an 'x there.

OpenStudy (waheguru):

yea sry it sould be 1

OpenStudy (abb0t):

Ok, now you have x = 0 and (2x-1) = 0 solve for "x" in the second equation by using ur algebra tools.

OpenStudy (waheguru):

(2x-1)=0 2x = 1 x=1/2

OpenStudy (waheguru):

Am I right?

OpenStudy (abb0t):

Yep. So your values for x = \(\frac{1}{2}\) and 0

OpenStudy (waheguru):

But how is the second value 0, that I do not understnad

OpenStudy (abb0t):

x(2x-1) = 0 ^ I I because you're setting these equal to zero. Hence x = 0 and x = (2x-1)